If the probability of a person catching a cold during cold and flu season is .4, 10 people are chosen at random, what is the probability that exactly four of them will catch a cold?
I know the answer but I have no idea how to solve the problem!
thanks!

Answer by sudhanshu_kmr(1152)

probability that a person catching a cold is .4 i.e P(A) = 4/10
= 2/5

now, probability that not catching a cold i.e P'(A)= 1-2/5 = 3/5

probability that exactly 4 person among 10 will catch a cold
= 10C4 * (2/5)^4 * (3/5)^6
= 210 * (16/625) * (729/15625)

try to get the concept...
u r welcome to get any help..